OpenStudy (anonymous):

What is the slope of the line passing through (0,3) and (4,5)?

OpenStudy (mertsj):

(5-3)/(4-0)=2/4=1/2

OpenStudy (anonymous):

thanks! can you answer this one too? http://openstudy.com/study#/updates/4f1dd440e4b04992dd245665

OpenStudy (mertsj):

-5

OpenStudy (mertsj):

y = -5x + 8 It's y = mx +b form and m = -5 so that is the slope.